OpenStudy (anonymous):

(k-6)(k+8) can someone help me i wanna know if i got the right answer

OpenStudy (anonymous):

it should be \[k ^{2}+2k-48\]

OpenStudy (anonymous):

yay i did it right :D

OpenStudy (anonymous):

good job! u got the concept :)

OpenStudy (anonymous):

yes thank u :)

hero (hero):

Did you do this: \((k-6)(k+8)\) \(\\= k(k + 8) - 6(k + 8) \\= k^2 + 8k - 6k - 48 \\= k^2 + 2k - 48\)

OpenStudy (anonymous):

yes i did :D & can you tell me what you got for this i wanna know if my answer is right (5m-2)(m+3)??

hero (hero):

Tell me what you got first, then I'll tell you if it is correct. Also show me what steps you did.

hero (hero):

@1108, please do not

OpenStudy (anonymous):

i wont i was gonna say you are right! lol

OpenStudy (anonymous):

\[5m ^{2}+13m-6\] i got that is it right ?

OpenStudy (anonymous):

yes mam

hero (hero):

Where are your steps @naomi_29

OpenStudy (anonymous):

but i got it right u really want me to show u all my steps i have more question to answer .-.

hero (hero):

Yes, I want to see you show your steps at least once so I know that you know.

OpenStudy (anonymous):

it doesnt let me put im going to take a picture of my work better k

hero (hero):

Okay....

OpenStudy (anonymous):

k there u cant really see it good

hero (hero):

Yeah, but you used FOIL method rather than Distribution method.

OpenStudy (anonymous):

okay thank u for the help :D